OpenStudy (anonymous):

4 square root of (x-3)-square root of(6x-17)=3

OpenStudy (anonymous):

\[4\sqrt{x-3}-\sqrt{6x-17}=3\]

OpenStudy (anonymous):

worked this down to quadratic form but wanted to see if im anywhere close

OpenStudy (anonymous):